sarah is 15 yrs older than joseph. in ten years from now. sarah will be twice as old as joseph .how old is sarah now?
Answer by Finavon(81)

Sarah S; Joseph J
Now S=J+15
Later (S+10)=2(J+10)
Substitute for S
J+15+10=2J + 20
Rearrange
5=J
So S=20
Sarah is 20 yrs